# Joint Channel and Data Estimation for Multiuser Extremely Large-Scale MIMO Systems

### v2

@@ I Introduction
To meet the demands for high spectral efficiency in future 6G systems [1], it is essential to further exploit spatial multiplexing and abundant spectral resources at mid/high frequency bands such as centimeter-wave (cmWave), milimeter-wave (mmWave), and sub-terahertz (sub-THz). In light of these requirements, extremely large-scale multiple-input-multiple-output (XL-MIMO) [2, 3, 4] has emerged as a promising technology, enabling sharp directive beamforming and extensive spatial multiplexing. However, the significant increase in antenna aperture leads to an expansion of the Rayleigh distance [5, 6], defined as the border between the near-field and far-field regions. Thus, the near-field effects in XL-MIMO systems may not be negligible in some practically-relevant circumstances, such as in small area coverage with high carrier frequency bands [7].
@@
Unlike the conventional far-field channel, the near-field channel depends not only on gains and angles (e.g., angles of arrivals (AoAs)) but also on distances from signal sources such as user equipments (UEs) and scatterers. Hence, conventional channel estimation methods such as [8, 9], which exploit the beam-domain sparsity under the assumption of planar wavefront, experience significant performance degradation in the near-field due to energy leakage effects in the beam-domain. To tackle this issue, the authors in [7] have proposed a polar-domain simultaneous orthogonal matching pursuit (P-SOMP) algorithm, which leverages the angle and distance sparsity known as polar-domain sparsity arising from the aforementioned near-field peculiar characteristics. In P-SOMP, polar (angle-distance) grids are generated by spatially quantizing the polar-domain to utilize compressed sensing techniques, however, in multiuser systems, P-SOMP requires orthogonal pilots to separate multiple UEs. As such, the proposed approach results in non-negligible overhead as the number of UEs grows, especially in XL-MIMO systems capable of spatially multiplexing many UEs.
@@
Considering these challenges, a near-field channel estimation algorithm, which works even with non-orthogonal pilots, has recently been proposed in [10] in the context of grant-free XL-MIMO systems [^1] Note that while this method was originally developed for jointly active user detection and channel estimation, it is also applicable to sole channel estimation problems without active user detection.. However, due to the non-orthogonality among pilots, inter-user interference still remains, so it is necessary to jointly estimate all UE channel components. As a result, this joint estimation significantly increases computational complexity because it requires UE-wise polar grids, which leads to a large grid size. Therefore, the authors in [10] proposed a 2D-compressive sampling matching pursuit (CoSaMP) algorithm, which is based on the CoSaMP algorithm in the polar-UE 2D domain constructed by UE-wise polar grids. While the 2D-CoSaMP algorithm can mitigate computational complexity, its estimation performance is hindered by overfitting to noisy measurements, which results from the inverse operation on over-sampled estimates. Consequently, subsequent data detection suffers from severe performance deterioration, particularly with high-order modulation.
@@
One of the prospective solutions to obtain accurate channel estimate with non-orthogonal pilots is joint channel and data estimation (JCDE) [11, 12, 13, 14, 15], where not only pilot sequences but also estimated data symbols are utilized as pilot replicas, leveraging their statistical quasi-orthogonality. The JCDE problem can be formulated as a bilinear inference problem (BIP). One of the prominent algorithms based on a Bayesian framework for BIP is bilinear generalized approximate message passing (BiGAMP) [16]. BiGAMP is an extension of GAMP, originally designed for a high-dimensional generalized-linear problem by utilizing loopy belief propagation (BP) with central limit theorem (CLT) and Taylor-series approximations based on large system limit to simplify the BP update. Due to the heavy dependency on the large system assumption of BiGAMP, the convergence performance deteriorates significantly in case the system is too small, the pilots are too short or when there is an improper prior distribution [11]. To address these issues, the authors in [11] have proposed bilinear Gaussian belief propagation (BiGaBP), which relaxes the BP update rules of BiGAMP, based on GaBP [17], without heavily relying on the approximation under a large system limit assumption. This relaxation of the approximation leads to performance improvements while maintaining the same complexity order as BiGAMP.
@@
Bilinear inference algorithms that exploit physical model structures, such as channel sparsity in the beam-domain, have been investigated in [13, 14]. In these papers, the channel sparsity is modeled using a Bernoulli Gaussian (BG) prior distribution because this prior is analytically tractable with a closed-form posterior. However, the sparse structure cannot be exactly expressed by the analytically tractable prior, which leads to modeling errors resulting in performance deterioration. To tackle this issue, the authors of [12] have integrated a model-based deterministic approach [18] into a Bayesian inference framework [11], referred to as AoA-aided BiGaBP. This deterministic approach rectifies the model mismatch caused by the use of the tractable prior.
@@
However, since this method assumes a far-field model, the model correction by the deterministic estimation is insufficient for the near-field region in XL-MIMO systems. Moreover, AoA-aided BiGaBP relies on a maximum ratio combining (MRC)-based detector, which cannot address the correlation caused by energy leakage in the beam-domain due to near-field effects. In addition, the computational complexity of the data denoising process in AoA-aided BiGaBP based on prior information for modulation constellations scales proportionally to not only the modulation order but also the number of antennas. This increase in complexity stems from the fact that BiGaBP suppresses the self-feedback of messages in the algorithmic iteration by generating antenna-wise extrinsic values based on BP rules without the Onsager correction term [11]. Consequently, the number of inputs to the denoiser function, based on modulation constellations, increases proportionally to the number of antennas.
@@
  • Initial channel estimation for JCDE: A novel initialization mechanism for the multiuser near-field channel estimation problem with pilot contamination due to non-orthogonal pilots is proposed, enabling an accurate initial estimate that is then used in the subsequent JCDE algorithm. The proposed initial channel estimation algorithm consists of two stages to maintain low computational complexity. In the first stage, angle and distance parameters for all UEs are estimated from the polar grids using the simultaneous orthogonal matching pursuit (SOMP) algorithm without pairing between each estimated path and the corresponding UE. Subsequently, the second stage involves the UE-path pairing using 2D-OMP [19] with a reduced number of grids constructed on the angle and distance parameters derived from the first stage. Owing to the above two-stage procedure, our proposed initial channel estimation outperforms the existing state-of-the-art scheme [10], while maintaining comparable computational complexity.
@@
  • JCDE algorithm with model-based estimation: A novel bilinear JCDE inference algorithm is proposed, which integrates a model-based deterministic estimation mechanism with a Bayesian inference framework to address possible performance degradation due to modeling errors in the prior knowledge assumed in the Bayesian inference, as in [11]. In contrast to the state-of-the-art AoA-aided BiGaBP under the assumption of far-field propagation, our algorithm estimates the channels as an aggregation of two distinct quantities: 1) a model-based estimate that captures the near-field channel structure and 2) its modeling error that captures how much different the current estimate is from the true channel. The model-based estimate is alternately updated through a matching pursuit algorithm exploiting the near-field model structures, whereas the residual modeling errors and data symbols are jointly estimated by the expectation propagation (EP) algorithm [20], where an approximate posterior is calculated by minimizing the Kullback-Leibler (KL) divergence between the true posterior and the approximate posterior. To tackle the spatial correlation caused by the energy leakage across neighboring beams in the beam domain while reducing the complexity, we introduce a novel posterior calculation design that enables the implementation of a sub-array-wise linear minimum mean square error (LMMSE)-based filter, allowing parallel computation of the matrix inversion with a much smaller dimension than the array size. This design indeed results in lower computational complexity compared to the state-of-the-art method, owing to the modification of an extrinsic value generation that does not rely on BP rules, as shown in the simulation results.
@@
Notation: The notation $[\mathbf{A}]_{i,j}$ indicates the $(i,j)$ element of the matrix $\mathbf{A}$. For a random variable $x$ and a probabilistic density function $p(x)$, $\mathbb{E}_{p(x)}[x]$ indicate the expectation of $x$ over $p(x)$. For any function $f(\mathbf{z})$, $\int_{\mathbf{z}/z_{i}}f(\mathbf{z})$ denotes the integral of $f(\mathbf{z})$ with respect to $\mathbf{z}$ except for $z_{i}$. The operator $\otimes$ denotes the Kronecker product. For the index sets =$\mathcal{I}=\{1,2,\ldots,I\}$ and $\mathcal{J}=\{1,2,\ldots,J\}$, ×$\mathcal{I}\times\mathcal{J}$ denotes the cartesian product of $\mathcal{I}$ and $\mathcal{J}$. $\mathcal{I}\setminus i$ represent the set $\{1,\ldots,i-1,i+1,\ldots,I\}$.
@@ II System Model
Figure 1: Near-field channel model.
Figure 2: The amplitude of the channel vector in the beam-domain, $\mathbf{h}_{u}$ for $N=200,L=1,f_{\mathrm{c}}=200\ \mathrm{GHz}$. One UE is located $r$ meters away from the BS ($r=\infty$ corresponds to the far-field case).
@@ II-A Channel Model
The near-field channel in the spatial domain between a BS and the $u$-th UE is modeled as
@@
where $\theta_{u,l}\in\mathbb{R}$, and $z_{u,l}\in\mathbb{C}$ denote the AoA and path gain of the $l$-th path and the $u$-th UE, respectively [21].
Without loss of generality, $l=1$ represents the line-of-sight (LoS) component and $l\in\{2,\ldots\hat{L}_{u}\}$ represents the non-line-of-sight (NLoS) components. Let $L\triangleq\sum_{u=1}^{U}L_{u}$ denote the total number of path including $U$-UEs. Accordingly, $r_{u,1}\in\mathbb{R}$ denotes the distance between the BS and the $u$-th UE, and $r_{u,l}\in\mathbb{R},\ l\neq 1$ is the distance between the BS and the $l$-th scatterer around the $u$-th UE. Besides, $\mathbf{a}(\theta_{u,l},r_{u,l})\in\mathbb{C}^{N\times 1}$ denotes the array response vector defined as
@@
where $r_{u,l}^{(n)}=\sqrt{r_{u,l}^{2}+y^{(n)2}-2r_{u,l}y^{(n)}\sin\theta_{u,l}}$ is the distance between the $n$-th antenna and the $u$-th UE or scatterers.
@@
The array response vector in the far-field region, i.e., when $r_{u,l}\rightarrow\infty$ is expressed as $[\mathbf{a}(\theta_{u,l},\infty)]_{n}=\exp\left[j\frac{2\pi y^{(n)}}{\lambda}\sin\theta_{u,l}\right]$ from (2). As the far-field array response depends only on the angle, the far-field channel $\mathbf{h}_{u}^{\mathcal{S}}$ can be converted into a sparse beam-domain channel $\mathbf{h}_{u}=\mathbf{D}_{N}\mathbf{h}_{u}^{\mathcal{S}}$ with the discrete Fourier transform (DFT) matrix $\mathbf{D}_{N}\in\mathbb{C}^{N\times N}$. In contrast, as the far-field approximation does not hold in the near-field region, the beam-domain near-field channel $\mathbf{h}_{u}$ exhibits not a simple sparse structure but rather a cluster sparse structure, which is caused by energy leakage due to a model mismatch between the DFT matrix $\mathbf{D}_{N}$ and the near-field array response $\mathbf{a}(\theta_{u,l},r_{u,l})$. To illustrate the energy leakage effects, Fig. 2 depicts the amplitude of the beam-domain channel vector in the near-field and far-field regions. It can be seen that the far-field channel possesses a distinct sparse structure with a peaky spike. On the other hand, the near-field channel exhibits a clustered sparsity with flatter peaks due to energy leakage. Hence, conventional channel estimation methods exploiting the beam-domain sparsity [8, 9, 18] encounter significant performance degradation in the near-field.
@@ II-B Received Signal Model
To estimate the near-field channel, the $u$-th UE transmits a non-orthogonal pilot sequence $\mathbf{x}_{\mathrm{p},u}\in\mathbb{C}^{K_{\mathrm{p}}\times 1}$ and data symbol $\mathbf{x}_{\mathrm{d},u}\in\mathbb{C}^{K_{\mathrm{d}}\times 1}$ subsequently, where $K_{\mathrm{p}}(<U)$ and $K_{\mathrm{d}}$ are the length of pilots and data symbols. Each entry of $\mathbf{x}_{\mathrm{d},u}$ is randomly generated from a $Q$-quadrature amplitude modulation (QAM) constellation $\mathcal{X}\triangleq\{\mathcal{X}_{1},\ldots,\mathcal{X}_{Q}\}$ with average symbol energy $E_{\mathrm{s}}$. Then, the received pilot $\mathbf{Y}_{\mathrm{p}}^{\mathcal{S}}\in\mathbb{C}^{N\times K_{\mathrm{p}}}$ and data $\mathbf{Y}_{\mathrm{d}}^{\mathcal{S}}\in\mathbb{C}^{N\times K_{\mathrm{d}}}$ in the spatial domain are given by
@@
where $\mathbf{X}_{\mathrm{p}}\triangleq\begin{bmatrix}\mathbf{x}_{\mathrm{p},1},\ldots,\mathbf{x}_{\mathrm{p},U}\end{bmatrix}^{\mathrm{T}}\in\mathbb{C}^{U\times K_{\mathrm{p}}}$ and $\mathbf{X}_{\mathrm{d}}\triangleq\begin{bmatrix}\mathbf{x}_{\mathrm{d},1},\ldots,\mathbf{x}_{\mathrm{d},U}\end{bmatrix}^{\mathrm{T}}\in\mathbb{C}^{U\times K_{\mathrm{d}}}$ are the transmitted pilot matrix and data matrix. $\mathbf{N}_{\mathrm{p}}\in\mathbb{C}^{N\times K_{\mathrm{p}}}$ and $\mathbf{N}_{\mathrm{d}}\in\mathbb{C}^{N\times K_{\mathrm{d}}}$ are the additive white Gaussian noise (AWGN) matrices, whose entries are generated from $\mathcal{CN}(0,\sigma^{2})$ with noise variance $\sigma^{2}$. By stacking the received pilot $\mathbf{Y}_{\mathrm{p}}^{\mathcal{S}}$ and data $\mathbf{Y}_{\mathrm{d}}^{\mathcal{S}}$, the effective received signal becomes $\mathbf{Y}^{\mathcal{S}}\triangleq\begin{bmatrix}\mathbf{Y}_{\mathrm{p}}^{\mathcal{S}},\mathbf{Y}_{\mathrm{d}}^{\mathcal{S}}\end{bmatrix}\in\mathbb{C}^{N\times K}$, and the sum length of pilots and data $K\triangleq K_{\mathrm{p}}+K_{\mathrm{d}}$, is formulated as

### v3

@@ I Introduction
Extremely large-scale multiple-input-multiple-output (XL-MIMO) has emerged as a promising technology, enabling sharp directive beamforming and extensive spatial multiplexing [2, 3]. However, the significant increase in antenna aperture leads to an expansion of the Rayleigh distance [4, 5], defined as the border between the near-field and far-field regions. Thus, the near-field effects in XL-MIMO systems may not be negligible in some practically-relevant circumstances, such as in small area coverage with high carrier frequency [6].
@@
Unlike the conventional far-field, the near-field channel depends not only on angles but also on distances from signal sources such as user equipments (UEs) and scatterers. Hence, conventional channel estimation methods such as [7, 8], which exploit the beam-domain sparsity under the assumption of planar wavefront, experience performance degradation in the near-field due to energy leakage effects in the beam-domain. To tackle this issue, the authors in [6] have proposed a polar-domain simultaneous orthogonal matching pursuit (P-SOMP) algorithm, which leverages the angle and distance sparsity known as polar-sparsity arising from the near-field peculiar characteristics. In P-SOMP, polar (angle-distance) grids are generated by spatially quantizing the polar-domain to utilize compressed sensing techniques. However, in multiuser systems, P-SOMP requires orthogonal pilots to separate multiple UEs. As such, this approach results in non-negligible overhead as the number of UEs grows, especially in XL-MIMO systems capable of spatially multiplexing many UEs.
@@
Considering these challenges, a near-field channel estimation algorithm, which works even with non-orthogonal pilots, has been proposed in [9]. However, due to the non-orthogonality among pilots, inter-user interference still remains, so it is necessary to jointly estimate all UE channel components. As a result, this joint estimation significantly increases computational complexity because it requires UE-wise polar grids, which leads to a large grid size. Therefore, the authors in [9] proposed a 2D-compressive sampling matching pursuit (CoSaMP) algorithm, which is based on the CoSaMP algorithm in the polar-UE 2D domain constructed by UE-wise polar grids. While the 2D-CoSaMP algorithm can mitigate computational complexity, its estimation performance is hindered by overfitting to noisy measurements, which results from the inverse operation on over-sampled estimates. Consequently, subsequent data detection suffers from severe deterioration, particularly with high-order modulation.
@@
One of the prospective solutions to obtain accurate channel estimate with non-orthogonal pilots is joint channel and data estimation (JCDE) [10, 11, 12, 13, 14], where not only pilot sequences but also estimated data symbols are utilized as pilot replicas. The JCDE problem can be formulated as a bilinear inference problem (BIP). One of the prominent algorithms based on a Bayesian framework for BIP is bilinear generalized approximate message passing (BiGAMP) [15]. BiGAMP is an extension of GAMP, originally designed for a high-dimensional generalized-linear problem by utilizing loopy belief propagation (BP) with central limit theorem (CLT) and Taylor-series approximations based on large system limit to simplify the BP update111 Assuming a bilinear inference problem from the noisy measurement $\mathbf{Y}=\mathbf{H}\mathbf{X}+\mathbf{N}$ with objective matrices $\mathbf{H}\in\mathbb{C}^{N\times U}$ and $\mathbf{X}\in\mathbb{C}^{U\times K}$, and a noise matrix $\mathbf{N}\in\mathbb{C}^{N\times K}$, the large system limit means that the dimensions $N$, $U$, and $K$ tend to infinity while the rates $N/U$ and $K/U$ remain constant [15]. . Due to the heavy dependency on the large system assumption of BiGAMP, the convergence performance deteriorates significantly when the system size is insufficient, the pilot length is short, or the prior distribution is misspecified [10]. To address these issues, the authors in [10] have proposed bilinear Gaussian belief propagation (BiGaBP), which relaxes the BP update rules of BiGAMP, based on GaBP [16], without heavily relying on the approximation under a large system limit assumption. This relaxation of the approximation leads to performance improvements while maintaining the same complexity order as BiGAMP.
@@
Bilinear inference algorithms that exploit physical model structures, such as channel sparsity in the beam-domain, have been investigated in [12, 13]. In these papers, the channel sparsity is modeled using a Bernoulli Gaussian (BG) prior distribution because this prior is analytically tractable with a closed-form posterior. However, the sparse structure cannot be exactly expressed by the analytically tractable prior, which leads to modeling errors. To tackle this issue, the authors of [11] have integrated a model-based deterministic approach [17] into a Bayesian inference framework [10], referred to as AoA-aided BiGaBP. This deterministic approach rectifies the model mismatch caused by the use of the tractable prior.
@@
However, since this method assumes a far-field model, the model correction via the deterministic approach is insufficient in the near-field, where the channel in the beam-domain exhibits cluster sparsity instead of simple sparsity with a sharp peak, due to energy leakage in the beam-domain. Moreover, AoA-aided BiGaBP relies on a maximum ratio combining (MRC)-based detector, which cannot address the correlation caused by the energy leakage. While the computational complexity of the MRC-based detection is relatively low, the data denoising process in AoA-aided BiGaBP, which relies on prior information about modulation constellations, requires significant computational complexity that scales proportionally not only to the modulation order but also to the number of antennas due to its reliance on the BP update rules. Hence, a channel estimation with a deterministic approach tailored to near-field structure, along with a computationally efficient data detection to address the energy leakage, is essential for the JCDE in the near-field region. Furthermore, an accurate initial channel estimation is crucial for JCDE algorithms because the performance of data and channel estimation heavily depends on the accuracy of the initial estimate, where an inaccurate initial estimate might lead to error propagation in the JCDE iterations.
@@
  • Initial channel estimation for JCDE: A novel initialization mechanism for the multiuser near-field channel estimation problem with pilot contamination due to non-orthogonal pilots is proposed, enabling an accurate initial estimate that is then used in the subsequent JCDE algorithm as the starting values. The proposed initial channel estimation algorithm consists of two stages to maintain low computational complexity. In the first stage, angle and distance parameters for all UEs are estimated from the polar grids using the simultaneous orthogonal matching pursuit (SOMP) algorithm without pairing between each estimated path and the corresponding UE. Subsequently, the second stage involves the UE-path pairing using 2D-OMP [18] with a reduced number of grids constructed on the angle and distance parameters derived from the first stage. Owing to the above two-stage procedure, our proposed initial channel estimation outperforms the existing state-of-the-art scheme [9], while maintaining comparable computational complexity.
@@
  • JCDE algorithm with model-based estimation: A novel bilinear JCDE inference algorithm is proposed, which integrates a model-based deterministic estimation mechanism with a Bayesian inference to address the modeling errors in a prior, as in [10]. In contrast to the state-of-the-art AoA-aided BiGaBP under the assumption of far-field, our algorithm estimates the channels as an aggregation of two distinct quantities: 1) a model-based estimate that captures the near-field channel structure and 2) its modeling error that captures how different the current estimate is from the true channel. The model-based estimate is alternately updated through a matching pursuit algorithm exploiting the near-field model structures, whereas the residual modeling errors and data symbols are jointly estimated by the expectation propagation (EP) algorithm [19], where an approximate posterior is calculated by minimizing the Kullback-Leibler (KL) divergence. To tackle the correlation caused by the energy leakage in the beam-domain while reducing the complexity, we introduce a novel posterior calculation design that enables the implementation of a sub-array-wise linear minimum mean square error (LMMSE)-based filter, allowing parallel computation of the matrix inversion with a smaller dimension than the array size. This design results in lower computational complexity compared to the state-of-the-art method, owing to the modification of an extrinsic value generation that does not rely on BP rules.
@@
Notation: The notation $\mathbf{A}]_{i,j}$ indicates the $(i,j)$ element of the matrix $\mathbf{A}$. $(\cdot)^{*}$, $(\cdot)^{\mathrm{T}}$, and $(\cdot)^{\mathrm{H}}$ indicate conjugate, transpose, and conjugate transpose, respectively. A block diagonal matrix consisting of matrices $\mathbf{A}_{1},\ldots\mathbf{A}_{N}$ is represented as $\mathrm{blkdiag}(\mathbf{A}_{1},\ldots,\mathbf{A}_{N})$. For a random variable $x$ and a probabilistic density function $p(x)$, $\mathbb{E}_{p(x)}[x]$ indicate the expectation of $x$ over $p(x)$. For any function $f(\mathbf{z})$, $\int_{\mathbf{z}/z_{i}}f(\mathbf{z})$ denotes the integral of $f(\mathbf{z})$ with respect to $\mathbf{z}$ except for $z_{i}$. The operator $\otimes$ denotes the Kronecker product. For the index sets $\mathcal{I}=\{1,2,\ldots,I\}$ and $\mathcal{J}=\{1,2,\ldots,J\}$, $\mathcal{I}\times\mathcal{J}$ denotes the cartesian product of $\mathcal{I}$ and $\mathcal{J}$. $\mathcal{I}\setminus i$ represent the set $\{1,\ldots,i-1,i+1,\ldots,I\}$. The notation $x\in[x_{1},x_{2}]$ denotes that $x$ belongs to the closed interval between $x_{1}$ and $x_{2}$.
@@ II System Model
Figure 1: Near-field channel model. $r_{u,l}^{(n)}$ is the distance between the $n$-th antenna and the $u$-th UE ($l=1$) or scatterers ($l\neq 1$).
Figure 2: The amplitude of the channel vector in the beam-domain, $\mathbf{h}_{u}$ for $N=200,L=1,f_{\mathrm{c}}=100\ \mathrm{GHz}$. One UE is located $r$ meters away from the BS ($r=\infty$ corresponds to the far-field case).
@@ II-A Channel Model
The near-field channel in the spatial-domain between the BS and the $u$-th UE is modeled as
@@
where $\theta_{u,l}\in\mathbb{R}$, and $z_{u,l}\in\mathbb{C}$ denote the angles of arrival (AoA) and path gain of the $l$-th path and the $u$-th UE, and $L_{u}$ is the number of paths for the $u$-th UE [20].
Without loss of generality, $l=1$ represents the line-of-sight (LoS) component and $l\in\{2,\ldots L_{u}\}$ represents the non-line-of-sight (NLoS) components. Let $L\triangleq\sum_{u=1}^{U}L_{u}$ denote the total number of paths including $U$-UEs. Accordingly, $r_{u,1}\in\mathbb{R}$ denotes the distance between the BS and the $u$-th UE, and $r_{u,l}\in\mathbb{R},\ l\neq 1$ is the distance between the BS and the $l$-th scatterer around the $u$-th UE. Besides, $\mathbf{a}(\theta_{u,l},r_{u,l})\in\mathbb{C}^{N\times 1}$ denotes the array response vector defined as
@@
where $r_{u,l}^{(n)}=\sqrt{r_{u,l}^{2}+y^{(n)2}-2r_{u,l}y^{(n)}\sin\theta_{u,l}}$ is the distance between the $n$-th antenna and the $u$-th UE ($l=1$) or scatterers ($l\neq 1$).
@@
The array response vector in the far-field region, i.e., when $r_{u,l}\rightarrow\infty$ is expressed as $\mathbf{a}(\theta_{u,l},\infty)]_{n}=\exp\left[j\frac{2\pi y^{(n)}}{\lambda}\sin\theta_{u,l}\right]$ from (2). As the far-field array response depends only on the angle, the far-field channel $\mathbf{h}_{u}^{\mathcal{S}}$ can be converted into a sparse beam-domain channel $\mathbf{h}_{u}=\mathbf{D}_{N}\mathbf{h}_{u}^{\mathcal{S}}$ with the discrete Fourier transform (DFT) matrix $\mathbf{D}_{N}\in\mathbb{C}^{N\times N}$. In contrast, as the far-field approximation does not hold in the near-field region, the beam-domain near-field channel $\mathbf{h}_{u}$ exhibits not a simple sparse structure but rather a cluster sparse structure, which is caused by energy leakage due to a model mismatch between the DFT matrix $\mathbf{D}_{N}$ and the near-field array response $\mathbf{a}(\theta_{u,l},r_{u,l})$. To illustrate the energy leakage effects, Fig. 2 depicts the amplitude of the beam-domain channel vector in the near-field and far-field regions. It can be seen that the far-field channel possesses a distinct sparse structure with a peaky spike. On the other hand, the near-field channel exhibits a clustered sparsity with flatter peaks due to energy leakage. Hence, conventional channel estimation methods exploiting the beam-domain sparsity [7, 8, 17] encounter significant performance degradation in the near-field.
@@ II-B Received Signal Model
To estimate the near-field channel, the $u$-th UE transmits a pilot sequence $\mathbf{x}_{\mathrm{p},u}\in\mathbb{C}^{K_{\mathrm{p}}\times 1}$ and data symbol $\mathbf{x}_{\mathrm{d},u}\in\mathbb{C}^{K_{\mathrm{d}}\times 1}$ subsequently, where $K_{\mathrm{p}}$ and $K_{\mathrm{d}}$ are the length of pilots and data symbols222 Since the pilots and data symbols are transmitted in distinct time symbols, unlike superimposed pilot transmission schemes [21], contamination between the pilots and data symbols does not occur. . To reduce the pilot overhead for channel estimation, we employ non-orthogonal pilots that satisfy $K_{\mathrm{p}}<U$ and $\mathbf{x}_{\mathrm{p},u}^{\mathrm{T}}\mathbf{x}_{\mathrm{p},u^{\prime}}^{\ast}\neq 0,(u^{\prime}\neq u)$, which results in the pilot contamination between UEs unlike the orthogonal pilots.
Each entry of $\mathbf{x}_{\mathrm{d},u}$ is randomly generated from a $Q$-quadrature amplitude modulation (QAM) constellation $\mathcal{X}\triangleq\{\mathcal{X}_{1},\ldots,\mathcal{X}_{Q}\}$ with average symbol energy $E_{\mathrm{s}}$. Then, the received pilot $\mathbf{Y}_{\mathrm{p}}^{\mathcal{S}}\in\mathbb{C}^{N\times K_{\mathrm{p}}}$ and data $\mathbf{Y}_{\mathrm{d}}^{\mathcal{S}}\in\mathbb{C}^{N\times K_{\mathrm{d}}}$ in the spatial-domain are given by
@@
where $\mathbf{X}_{\mathrm{p}}\triangleq\begin{bmatrix}\mathbf{x}_{\mathrm{p},1},\ldots,\mathbf{x}_{\mathrm{p},U}\end{bmatrix}^{\mathrm{T}}\in\mathbb{C}^{U\times K_{\mathrm{p}}}$ and $\mathbf{X}_{\mathrm{d}}\triangleq\begin{bmatrix}\mathbf{x}_{\mathrm{d},1},\ldots,\mathbf{x}_{\mathrm{d},U}\end{bmatrix}^{\mathrm{T}}\in\mathbb{C}^{U\times K_{\mathrm{d}}}$ are the transmitted pilot matrix and data matrices. $\mathbf{N}_{\mathrm{p}}\in\mathbb{C}^{N\times K_{\mathrm{p}}}$ and $\mathbf{N}_{\mathrm{d}}\in\mathbb{C}^{N\times K_{\mathrm{d}}}$ are the additive white Gaussian noise (AWGN) matrices, whose entries are generated from $\mathcal{CN}(0,\sigma^{2})$ with noise variance $\sigma^{2}$. By stacking the received pilot $\mathbf{Y}_{\mathrm{p}}^{\mathcal{S}}$ and data $\mathbf{Y}_{\mathrm{d}}^{\mathcal{S}}$, the effective received signal $\mathbf{Y}^{\mathcal{S}}\triangleq\begin{bmatrix}\mathbf{Y}_{\mathrm{p}}^{\mathcal{S}},\mathbf{Y}_{\mathrm{d}}^{\mathcal{S}}\end{bmatrix}\in\mathbb{C}^{N\times K}$, with $K\triangleq K_{\mathrm{p}}+K_{\mathrm{d}}$, is formulated as